### v1

# Cauchy activation function and XNet

Abstract
We have developed a novel activation function, named the Cauchy Activation Function. This function is derived from the Cauchy Integral Theorem in complex analysis and is specifically tailored for problems requiring high precision. This innovation has led to the creation of a new class of neural networks, which we call (Comple)XNet, or simply XNet.
We will demonstrate that XNet is particularly effective for high-dimensional challenges such as image classification and solving Partial Differential Equations (PDEs). Our evaluations show that XNet significantly outperforms established benchmarks like MNIST and CIFAR-10 in computer vision, and offers substantial advantages over Physics-Informed Neural Networks (PINNs) in both low-dimensional and high-dimensional PDE scenarios.
1 Introduction
In today’s scientific exploration, the rise of computational technology has marked a significant turning point. Traditional methods of theory and experimentation are now complemented by advanced computational tools that tackle the complexity of real-world systems. Machine learning, particularly deep neural networks, has led to breakthroughs in fields like image processing and language understanding [3, 7], and its application to scientific problems–such as predicting protein structures [9, 10] or forecasting weather [13]–demonstrates its potential to revolutionize our approach.
One of the primary challenges in computational mathematics and artificial intelligence (AI) lies in determining the most appropriate function to accurately model a given dataset. In machine learning, the objective is to leverage such functions for predictive purposes. Traditional methods rely on predetermined classes of functions, such as polynomials or Fourier series, which, though simple and computationally manageable, may limit the flexibility and accuracy of the fit. In contrast, modern deep learning neural networks primarily employ locally linear functions with nonlinear activations.
1.1 Algorithm Development
In our previous work [1], we introduced the initial concept of extending real-valued functions into the complex domain, using the Cauchy integral formula to device a machine learning algorithm, and demonstrating its high accuracy and performance through examples in time series. In this work, we introduce a more general method stemming from the same mathematical principles. This approach is not confined to addressing mathematical physics problems such as low-dimensional PDEs; it also effectively tackles a broad spectrum of AI application issues, including image processing. This paper primarily showcases examples in image processing and both low-dimensional and high-dimensional (100-dimensional) PDEs. We are actively continuing our research to explore further applications, with very promising preliminary results.
Recent advancements in deep learning for solving PDEs and CV capabilities are well-documented, with significant contributions from transformative architectures that integrate neural networks with PDEs to enhance function approximation in high-dimensional spaces [21, 22, 23, 24, 25]. In CV, comprehensive surveys and innovative methods have significantly advanced visual processing techniques [26, 27, 28, 29].
Also, we have extensively reviewed literature on integrating complex domains into neural network architectures. Explorations of complex-valued neural networks and further developments highlight the potential of complex domain methodologies in modern neural network frameworks [2, 11, 12, 31, 38, 39].
Feedforward Neural Networks (FNNs), despite their capabilities, are often limited by the granularity of approximation they can achieve due to traditional activation functions such as ReLU or Sigmoid [32, 33, 4]. Recent works aim to optimize network performance through the discovery of more effective activation functions, leading to significant advancements in network functionality and computational efficiency [40, 14, 15, 16, 17, 18, 19, 30].
We propose a novel method that does not directly incorporate complex numbers. Drawing on insights from the complex Cauchy integral formula, we utilize Cauchy kernels as basis functions and introduce a new Cauchy activation function. This innovative approach significantly enhances the network’s precision and efficacy across various tasks, particularly in high-dimensional spaces.
Figure 1: ReLU
Figure 2: Sigmoid
The Cauchy activation function can be expressed as follows
$\phi_{\lambda_{1},\lambda_{2},d}(x)=\frac{\lambda_{1}*x}{x^{2}+d^{2}}+\frac{\lambda_{2}}{x^{2}+d^{2}},$
where $\lambda_{1},\lambda_{2},d$ are trainable parameters.
Theoretically, the Cauchy activation function can approximate any smooth function to its highest possible order. Moreover, from the figure below, we observe that our activation function is localized and decays at the both ends. This feature turned out to be very useful in approximating local data. This capability to finely tune to specific data segments sets it apart significantly from traditional activation functions such as ReLU.
Figure 3: 2 terms of Cauchy activation
In the next section, we will explain how the Cauchy activation function is derived and why it is mathematically efficient.
2 Approximation Theorems
Consider, for example, a dataset comprising values of certain function $g(x_{i}),i=1,\ldots,n$ corresponding to points $x_{1},\ldots,x_{n}$ on the real line. These data may contain noises. We begin by assuming that the target function for fitting is a real-analytic function $f$. Although this assumption may appear stringent, it is important to note that non-analytic functions, due to their unpredictable nature and form, exhibit a weak dependence on specific data sets. Another perspective is that our aim is to identify the most suitable analytic function that best fits the provided dataset.
Real analytic functions can be extended to complex plane. The central idea of our algorithm is to place observers in the complex plane. Similar to activation for each node in artificial neural network, a weight is computed and assigned to each observer. The value of the predicted function $f$ at any point is then set to be certain weighted average of all observers. Our core mathematical theory is the Cauchy Approximation Theorem (Section 2, Theorem 1), derived in the next section. The Cauchy Approximation Theorem guarantees the efficacy and the accuracy of the predicted function. Comparing with the Universal Approximation Theorem for artificial neural network, whose proof takes considerable effort, our theorem comes directly from Cauchy Integration formula (eq 1).
In section 2, we start with Cauchy integral for complex analytical functions, leading us to the derivation the Cauchy Approximation Theorem. This theorem serves as the mathematical fundation of our algorithm. Theoretically, our algorithm can achieve a convergence rate of arbitrarily high order in any space dimensions.
In Sections 3, we evaluate our algorithm using a series of test cases. Observers are manually positioned within the space, and some known functions are employed to generate datasets in both one-dimensional and two-dimensional spaces. The predicted functions are then compared with the actual functions. As anticipated, the results are exceptional. We also conducted tests on datasets containing random noises, and the outcomes were quite satisfactory. The algorithm demonstrates impressive predictive capabilities when it processes half-sided data to generate a complete function.
We anticipate that this innovative algorithm will find extensive applications in fields such as computational mathematics, machine learning, and artificial intelligence. This paper provides a fundamental principle for the algorithm.
We will formally state the fundamental theory behind our algorithm.
Given a function $f$ holomorphic in a compact domain $U\subset\mathbb{C}^{N}$ in the complex $N$ dimensional space. For simplicity, we assume that $U$ has a product structure, i.e., $U=U_{1}\times U_{2}\times\ldots\times U_{n}$, where each $U_{k}$, $k=1,\ldots,N$ is a compact domain in the complex plane. Let $P$ be the surface defined by
$P=\partial U_{1}\times\partial U_{2}\times\ldots\times\partial U_{N}$
Then a multi-dimensional Cauchy integral formula for $U$ is given by:
[…]
for all $(z_{1},\ldots,z_{N})\in U$.
The magic of the Cauchy Integration Formula lies in its ability to determine the value of a function at any point using known values of the function. This concept is remarkably similar to the principles of machine learning!
The Cauchy integral can be approximated, to any prescribed precision, by a Riemann sum over a finite number of points on $P$. We can simplify the resulting Riemann sum in the following form:
$f(z_{1},z_{2},\ldots z_{N})\approx\sum_{k=1}^{m}\frac{\lambda_{k}}{(\xi_{1}^{k}-z_{1})(\xi_{2}^{k}-z_{2})\cdots(\xi_{N}^{k}-z_{N})},$ (2)
where $\lambda_{1},\lambda_{2},\ldots,\lambda_{m}$ are parameters depending on the functional values at the sample points on $P$.
The Cauchy integral Formula guarantees the accuracy of the above approximation if enough points on $P$ are taken. However, there is no reason at all that the sample points has to be on $P$ to achieve the best approximation. Indeed, the surface $P$ itself is quite arbitrary.
We can state our fundamental theorem.
Theorem 1 (Cauchy Approximation Theorem).
Let $f(z_{1},z_{2},\ldots z_{N})$ be an analytic function in an open domain $U\subset\mathbb{C}^{N}$ and let $M\subset U$ be a compact subset in $U$. Given any $\epsilon>0$, there is a list of points $(\xi_{1}^{k},\ldots,\xi_{N}^{k})$, for $k=1,2,\ldots,m$, in $U$ and corresponding parameters $\lambda_{1},\lambda_{2},\ldots,\lambda_{m}$, such that
$\vline\,f(z_{1},z_{2},\ldots z_{N})-\sum_{k=1}^{m}\frac{\lambda_{k}}{(\xi_{1}^{k}-z_{1})(\xi_{2}^{k}-z_{2})\cdots(\xi_{N}^{k}-z_{N})}\,\vline<\epsilon,$ (3)
for all points $(z_{1},z_{2},\ldots z_{N})\in M$.
As we have explained, the proof is a simple application of the Cauchy Theorem. We omit the details of the proof. We remark that, for any given $\epsilon>0$, the number of points $m$ needed is approximately at the level of $m\sim\epsilon^{-N}$, or equivalently, the error is approximately at $\epsilon\sim m^{-1/N}$. In fact, due to the nature of complex analytic functions, where the size of the function also bounds the derivative of the function, the error is much smaller. In fact, for any fixed integer $k>0$, one can show in theory that $\epsilon\sim o(m^{-k})$ for large $m$.
The Cauchy approximation is stated for complex analytical functions. If the original function is real, it can be simplified as follows.
$f(x_{1},x_{2},\ldots x_{N})\approx\mbox{Re}\left(\sum_{k=1}^{m}\frac{\lambda_{k}}{(\xi_{1}^{k}-x_{1})(\xi_{2}^{k}-x_{2})\cdots(\xi_{N}^{k}-x_{N})}\right),$ (4)
where $x_{1},x_{2},\ldots,x_{N}$ are real, $\lambda_{k},\;k=1,\ldots,m$ and $\xi_{i}^{j},\;i=1,\ldots N,\;j=1,\ldots,m$ are complex numbers with positive imaginary parts, Re($z$) is the real part of $z$.
A special simple case is when $N=1$, a function of one complex or real variable. In this case, the approximation is
$f(x)\approx\mbox{Re}\left(\sum_{k=1}^{m}\frac{\lambda_{k}}{\xi_{k}-x}\right),$ (5)
with $x$ real and $\xi_{k},\lambda_{k},\;k=1,\ldots m$ complex.
The Cauchy activation function can be easily derived from the above formula, by expressing all complex numbers in real and imaginary parts,
$\phi_{\lambda_{1},\lambda_{2},d}(x)=\frac{\lambda_{1}*x}{x^{2}+d^{2}}+\frac{\lambda_{2}}{x^{2}+d^{2}},$
where $\lambda_{1},\lambda_{2},d$ are all trainable real numbers.
Cauchy approximation theorem can be easily implemented into an artificial neural network ([1]). The resulting network, termed CauchyNet, is very efficient for lower dimensional problems, say for $N\leq 10$. However, for large $m$, the multiplicative terms in the denominator pose serious computational difficulties. As typical computer vision problems are high dimensional, with the input data for a $30\times 30$ pixel image being at least 900 dimensional, we need a different algorithm to handle high dimensional problems.
Towards this end, we prove a general approximation theorem. The central idea is that, if one can approximate one-dimensional function with certain functional class, then one can extend the functional class to approximate functions in any dimension. Mathematically, this corresponds to approximation in dual spaces. This method is particularly effective for feature-capturing.
As an example, Cauchy approximation can approximate any one dimensional function, as we see in the above theorem, therefore it can approximate any dimensioanl function through linear combinations.
More precisely, we consider $C(\mathbb{R})=C(\mathbb{R},\mathbb{R})$, the set of continuous real-valued functions in $\mathbb{R}$. A family of real valued functions on $\mathbb{R}$, $\Phi=\{\phi_{a}\;|\;a\in A\}$, where $A$ is some index set, is said to have universal approximation property or simply approximation property, if for any closed bounded interval $I$ in $\mathbb{R}$ and any continuous function $g\in C(I,\mathbb{R})$, there is a sequence of functions $g^{j}=a_{1}^{j}g_{a_{1}}+a_{2}^{j}g_{a_{2}}+\dots+a_{k_{j}}^{j}g_{a_{k_{j}}}$ such that $g^{j}\rightarrow g$ uniformly over $I$. In other words, $\Phi$ has approximation property if every continuous function can be approximated by a linear combination of functions in $\Phi$, uniformly over a bounded interval.
The main result is the following general approximation theorem.
Theorem 2 (General Approximation Theorem).
Let $\Phi$ be a family of functions in $C(\mathbb{R},\mathbb{R})$ with the universal approximation property. Let
$\Phi^{N}=\{\phi_{a}(a_{1}x_{1}+a_{2}x_{2}+\ldots+a_{N}x_{N})\;|\;(a_{1},\ldots,a_{N})\in\mathbb{R}^{N},\phi_{a}\in\Phi\}$
Then, the family of functions $\Phi^{N}$ has the universal approximation property in $C(\mathbb{R}^{N},\mathbb{R})$, i.e, every continuous function in $\mathbb{R}^{N}$ can be approximated by a linear combination of functions in $\Phi^{N}$, uniformly over a compact subset in $\mathbb{R}^{N}$.
The proof of this theorem can be found in Appendix A.
The Cauchy Approximation Theorem and the General Approximation Theorem demonstrate that the Cauchy activation function can be used effectively, at least mathematically, to approximate functions of any dimension. We will test its practical applications in the following sections.
Compared to popular neural networks, it appears that only the activation function needs to be changed. However, due to the novelty of the Cauchy activation function, structural changes to the neural networks are necessary to leverage its intrinsic efficiencies. Generally, we can significantly simplify complex networks to achieve comparable or better results. This theoretical and operational framework prompts us to rename the network to ’XNet’ to reflect the fundamental improvements in our networks capabilities.
3 Examples
In this section we focus on image classification and PDEs.
It is important to note that Theorem 2 applies to any higher orders. When solving PDEs, the higher-order accuracy of the solution is a key criterion for evaluating the method. As we optimize through parameter tuning, we will observe increasingly higher-order approximation effects. With the amount of data we have, we have already seen high-order accuracy. Since our theorem applies to arbitrary higher orders, we can observe even higher-order accuracy through parameter tuning and other techniques. This will be a focus of our future work.
3.1 MNIST
We try the XNet with a simple multi-scale kernel size convolution layer, and achieved the best acc o the testing data:
Handwriting recognition by a XNet. As shown in Fig. , the network consists of an input layer $W_{in}$, a hidden layer $W$ and an output layer $W_{out}$. The input $28\times 28$ grayscale image is reshaped into a $784\times 1$ vector and compressed by the input layer into $N$ features, which are then fed to the $N\times N$ hidden layer. The output layer maps the $N$ hidden outputs to 10 classes, representing digits from 0 to 9. The simulation model is built in and trained by ADAM Optimizer, with a learning rate 0.0001.
We used the simplest two-layer linear model with ReLU and Cauchy activation functions separately. It can be observed that with the same 20 epochs, the model using ReLU activation achieved $89.98\%$ accuracy on the validation set, while the model using the Cauchy activation function achieved $94.95\%$ accuracy. Additionally, after 20 epochs, the validation loss for the model using ReLU was 0.0019, whereas for the model using the Cauchy activation function, it was 0.0009.
Figure 4:
Figure 5:
Table 1: Validation Results for ReLU and Cauchy at lr=0.001
ReLU Cauchy Activation
Epochs Runtime (s) Valid Loss Valid Acc Runtime (s) Valid Loss Valid Acc
20 60 0.0019 90% 70 0.0009 95%
In the 2nd experiment, we used a simple convolutional neural network (CNN) with a single convolutional layer, followed by a max-pooling layer and a fully connected layer.
Figure 6: Training and validation performance of CNN with ReLU and Cauchy activation functions.
We experimented with two different activation functions: ReLU and Cauchy. In our model configuration, we replaced the ReLU activation function with the Cauchy activation function after the third convolutional layer. The learning rate was set to 0.001 for both models. The training and validation losses, as well as accuracies after 20 epochs, are presented in Table 2.
Table 2: Performance comparison of CNN with ReLU and Cauchy activation functions after 20 epochs, using a 3-layer convolutional architecture.
Activation LR Training Loss Training Acc Valid Loss Valid Acc
ReLU 0.001 0.0190 0.9935 0.0312 0.9905
Cauchy 0.001 0.0039 0.9988 0.0254 0.9924
As shown in Table 2, the model using the Cauchy activation function achieved a training loss of 0.0039 and a training accuracy of $99.88\%$, while the validation loss was 0.0254 and the validation accuracy was $99.24\%$. On the other hand, the model using the ReLU activation function resulted in a training loss of 0.0190 and a training accuracy of $99.35\%$, with a validation loss of 0.0312 and a validation accuracy of $99.05\%$.
These results suggest that the Cauchy activation model outperformed the ReLU model in both training and validation, showing lower losses and higher accuracies, indicating better generalization capabilities.
Remark We observed that placing the activation function with tunable parameters after the convolutional layer led to a rapid decrease in the loss. However, this approach also introduced a slight instability of validation accuracy during training.
3.2 CIFAR-10
In this study, we evaluated the performance of Cauchy activation function compared to the standard ReLU activation function, using the CIFAR-10 dataset. This dataset consists of 60,000 32x32 color images across 10 classes, making it a common benchmark for image classification tasks.
We employed two different model architectures:
  • CNN: A custom Convolutional Neural Network (CNN) designed to classify CIFAR-10 images. The architecture includes multiple convolutional layers followed by batch normalization and activation functions. Initially, ReLU was used in all layers, but in subsequent layers, ReLU was replaced with Cauchy activation, which is designed to capture more complex non-linear relationships using techniques from complex analysis.
  • ResNet9: A compact version of the ResNet architecture, known for its efficiency and effectiveness on smaller datasets like CIFAR-10. We modified the standard ResNet9 by replacing ReLU with Cauchy activation in the latter layers, aiming to improve the model’s ability to generalize.
Training Procedure Both models were trained in a similar fashion:
Cifar10CnnModel was trained in two phases: an initial training phase of 20 epochs with a learning rate of 0.001, followed by a fine-tuning phase of 10 epochs with a reduced learning rate of 0.0001. ResNet9 was trained for 8 epochs with a learning rate of 0.001. The standard ResNet9 model achieved a validation accuracy of $90\%$ after 8 epochs, while our model, which incorporated the Cauchy activation function, reached a higher accuracy of $92\%$ within the same number of epochs.
Table 3: Comparison of Model Performance on CIFAR-10
Model Convolutional Layers Training Epochs Accuracy
CNN (ReLU) 6 30 78.9%
CNN (Cauchy) 6 30 81.9%
ResNet9 (ReLU) 8 8 90.0%
ResNet9 (Cauchy) 8 8 92.0%
Figure 7: Validation acc: Comparison between Cauchy activation and ReLU in CNN
Figure 8: Validation acc: Comparison between Cauchy activation and ReLU in ResNet9 over 8 epochs.
Both models demonstrate that Cauchy activation can outperform ReLU in certain contexts, particularly when fine-tuning on complex datasets like CIFAR-10. The results indicate that Cauchy activation allows for faster convergence and higher final accuracy, making it a promising alternative to standard activation functions in deep learning models.
3.3 PDE: Heat Function
The 1-dimensional heat equation in this example is given by:
$\frac{\partial u}{\partial x}-2\frac{\partial u}{\partial t}-u=0,$ (6)
where $u(x,t)$ is the temperature distribution function, $x$ is the spatial coordinate, and $t$ is the time.
The boundary and initial conditions are specified as follows:
  • Initial condition:
    $u(x,0)=6e^{-3x},\quad\text{for}\;0\leq x\leq 2.$ (7)
  • Boundary conditions:
    $u(0,t)=0,\quad\text{for}\;0\leq t\leq 1,$ (8)
    $u(2,t)=0,\quad\text{for}\;0\leq t\leq 1.$ (9)
We trained the original PINN with the sigmoid activation function and then modified the network to use a Cauchy activation function instead of the sigmoid function. The training process was repeated with the new activation function to compare the performance.
Figure 9: loss curve by XNet and PINN
Table 4: Comparison with PINN
Activation Network Layers Training Loss Mean Error
Sigmoid 5 Layers FNN 0.0064 2e-3
Cauchy 5 Layers 0.0003 6e-5
3.4 Poisson Equation with Dirichelet Boundary Condition
We study the Poisson equation
$\nabla^{2}u(x,y)=f(x,y),\quad\text{for }(x,y)\in\Omega,$ (10)
with Dirichlet boundary conditions:
$u(x,y)=0,\quad\text{for }(x,y)\in\partial\Omega,$ (11)
where the source term $f(x,y)=-8\pi^{2}\sin(2\pi x)\sin(2\pi y)$. The ground truth solution is $u(x,y)={\rm sin}(2\pi x){\rm sin}(2\pi y)$.
The dataset size is 2000, 1000 interior points and 1000 boundary points.
In this simple low-dimensional equation, we firstly found that using the least-squares method in MATLAB effectively solves the problem. We worked with 1000 interior points and 1000 boundary points, placing 400 observation points, which correspond to the boundary points in the Cauchy integral formula in complex space. The method is equivalent with our CauchyNet in [1]. We defined the boundary as an ellipse. (If we optimize the boundary points, we can achieve even higher computational accuracy.) The results are shown below:
Figure 10: Cauchy activation
Figure 11: Analytic solution
Figure 12: Pointwise Difference between two solutions
The $L^{2}$ error is computed as follows: $\text{diff}=U_{\text{grid}}-F_{\text{grid}},$where $U_{\text{grid}}$ represents the predicted solution and $F_{\text{grid}}$ represents the analytical solution.
The $L^{2}$ error (mean squared error) is given by:
$L^{2}_{\text{error}}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(\text{diff}_{i})^{2}},$ (12)
where $n$ is the total number of grid points. The $L^{2}$ error is 0.0076886.
Next, we used a PINN model with a structure of size $[2,200,1]$, consisting of one hidden layer. The learning rate was set to 0.001 for the first 7000 epochs, and reduced to 0.0001 for the final 1000 epochs. The optimizer used was Adam. We compared the performance of two different activation functions: the tanh activation function and the Cauchy activation function over a total of 8000 training epochs.
Our loss function consists of two parts: the first part represents the Mean Squared Error (MSE) of the residuals from the equation, while the second part accounts for the MSE of the boundary conditions:
Loss $=\frac{1}{N_{\text{interior}}}\sum_{i=1}^{N_{\text{interior}}}\left(\frac{\partial^{2}u(x_{i})}{\partial x_{1}^{2}}+\frac{\partial^{2}u(x_{i})}{\partial x_{2}^{2}}-y_{\text{train},i}\right)^{2}+\frac{1}{N_{\text{boundary}}}\sum_{j=1}^{N_{\text{boundary}}}\left(u(x_{j})-u_{\text{boundary},j}\right)^{2}.$
Since the loss values were initially large and decreased significantly over time, the overall loss curve did not clearly highlight the differences between the methods. So we plotted the loss curve for the final 1000 epochs separately, where it becomes evident that our method (Cauchy activation) holds a clear advantage over the original PINN.
Figure 13: Loss curve for PINN and XNet
Activation Function Training loss
Tanh 0.0349
Cauchy 0.00354
Table 5: Comparison of activation functions and their respective training loss.
The results clearly demonstrate that our XNet model significantly outperforms the standard activation functions, as seen in the graph where the Cauchy activation function consistently achieves a lower loss, showing superior effectiveness.
In addition to the one-hidden-layer model, we also tested a PINN model with two hidden layers, with a structure of size $[2,20,20,1]$. The results for this two-hidden-layer model were similar to those of the one-hidden-layer model. Both models were trained using a learning rate of 0.001. After 8000 epochs, the tanh-based PINN achieved a training error of 0.119, while the XNet (Cauchy activation function) achieved a significantly lower training error of 0.0382.
3.4.1 High Dimensional PDE
In this section we test the XNet solver in the case of an 100-dimensional AllenCahn PDE with a cubic nonlinearity.
Assume for all $s,t\in[0,T]$, $x,w\in\mathbb{R}^{d}$, $y\in\mathbb{R}$, $z\in\mathbb{R}^{1\times d}$, $m\in\mathbb{N}$ that $\gamma_{m}=5\cdot 10^{-4}$, $d=100$, $T=\frac{3}{10}$, $N=20$, $\mu(t,x)=0$, $\sigma(t,x)w=\sqrt{2}w$, $\xi=(0,0,\ldots,0)\in\mathbb{R}^{d}$, $\Upsilon(s,t,x,w)=x+\sqrt{2}w$, $f(t,x,y,z)=y-y^{3}$, and $g(x)=\left[2+\frac{2}{5}\|x\|_{\mathbb{R}^{d}}^{2}\right]^{-1}$. Note that the solution $u$ satisfies for all $t\in[0,T)$, $x\in\mathbb{R}^{d}$ that $u(T,x)=g(x)$ and
$\frac{\partial u}{\partial t}(t,x)+u(t,x)-[u(t,x)]^{3}+\left(\Delta_{x}u\right)(t,x)=0.$
In our study, we employed a model identical to the one discussed in the paper [8]. We simplified the original model by reducing the multilayer perceptron (MLP) component to a single layer, effectively halving the parameter count. Then, we substituted the activation function and evaluated performance differences. We set up an Adam optimizer with a learning rate of 0.005. The comparison model configuration remained the same as described in the original paper [8] to ensure a fair evaluation.
Table 6: Comparison of Training Loss between Cauchy and ReLU
Step Cauchy Loss ReLU Loss Factor of Reduction
0 $1.5698\times 10^{-1}$ $1.5637\times 10^{-1}$ $\approx 1.00$
100 $2.9323\times 10^{-3}$ $9.8792\times 10^{-2}$ $\approx 33.70$
200 $3.3228\times 10^{-3}$ $7.8492\times 10^{-2}$ $\approx 23.63$
500 $2.9574\times 10^{-3}$ $3.7295\times 10^{-2}$ $\approx 12.61$
1000 $2.9667\times 10^{-3}$ $1.1308\times 10^{-2}$ $\approx 3.81$
2000 $2.3857\times 10^{-3}$ $4.9802\times 10^{-3}$ $\approx 2.09$
Figure 14: loss curve of Allen Cahn
Appendix A Proofs for Theorem 2
Proof: The proof is based on Stone-Weierstrauss Theorem: the set of all polynomials is dense in $C(B,\mathbb{R})$ for any compact subset $B\subset\mathbb{R}^{N}$. Since any polynomial is a linear combination of monomials, we just need to show that linear combinations of functions in $\Phi^{N}$ can approximate any monomial.
Since $\Phi$ has the approximation property, it therefore can approximate $x^{k}$ for any integer $k$, this implies that $\Phi^{N}$ can approximate any function of the form
$(a_{1}x_{1}+a_{2}x_{2}+\ldots+a_{N}x_{N})^{k},\;\mbox{ for }k\in\mathbb{N},\;(a_{1},\ldots,a_{N})\in\mathbb{R}^{N}.$
For any fixed integer $k$, the multinomial expansion of the above function is a linear combination of $C(k+N-1,N-1)$ monomial terms such as
$x_{1}^{k_{1}}x_{2}^{k_{2}}\cdots x_{N}^{k_{N}},\;\;k_{1}+k_{2}+\ldots,k_{N}=k.$
Here $C(k+N-1,N-1)$ is combinatorial number $k+N-1$ choosing $N-1$. We claim that the converse is also true, i.e., every such monomial can be written as a linear combination of functions of the form
$(a_{1}x_{1}+a_{2}x_{2}+\ldots+a_{N}x_{N})^{k}.$
Our theorem will directly follow from this claim.
We prove the claim by induction on $N$.
For $N=1$, this is obviously true.
For $N=2$,
$(x_{1}+0x_{2})^{k}$ $=$ $x_{1}^{k}$
$(x_{1}+1x_{2})^{k}$ $=$ $x_{1}^{k}+C(k,1)x_{1}^{k-1}x_{2}+\ldots+C(k,k-1)x_{1}x_{2}^{k-1}+x_{2}^{k}$
$(x_{1}+2x_{2})^{k}$ $=$ $x_{1}^{k}+C(k,1)2x_{1}^{k-1}x_{2}+\ldots+C(k,k-1)2^{k-1}x_{1}x_{2}^{k-1}+2^{k}x_{2}^{k}$
$(x_{1}+3x_{2})^{k}$ $=$ $x_{1}^{k}+C(k,1)3x_{1}^{k-1}x_{2}+\ldots+C(k,k-1)3^{k-1}x_{1}x_{2}^{k-1}+3^{k}x_{2}^{k}$
$(x_{1}+kx_{2})^{k}$ $=$ $x_{1}^{k}+C(k,1)kx_{1}^{k-1}x_{2}+\ldots+C(k,k-1)k^{k-1}x_{1}x_{2}^{k-1}+k^{k}x_{2}^{k}$
This is a system of equations for monomials of the form $x_{1}^{j}x_{2}^{k-j}$. The coefficient matrix is obviously nonsingular, therefore each $x_{1}^{j}x_{2}^{k-j}$ can be solved as a linear combination of $(x_{1}+jx_{2})^{k}$, with $j=0,1,\ldots,k$.
For $N=3$, fix any real number $l$, we can expand $((x_{1}+lx_{2})+jx_{3})^{k}$ for $j=0,1,\ldots,k$ into the product of the form
$(x_{1}+lx_{2})^{k-k_{3}}x_{3}^{k_{3}},\;\mbox{ for }k_{3}=1,...,k-1.$
Same as the previous case with $N=2$, each of the above terms can be expressed as a linear combination of
$((x_{1}+lx_{2})+jx_{3})^{k},\;\mbox{ for }\;j=0,1,\ldots,k.$
Now for any fixed $k_{3}$, we choose $l=0,1,\ldots,k-k_{3}$ and expand
$(x_{1}+lx_{2})^{k-k_{3}}x_{3}^{k_{3}}.$
Similar to the case with $N=2$, we now obtain $k-k_{3}$ independent equations. From these equations, we can solve each monomial
$x_{1}^{k_{1}}x_{2}^{k_{2}}x_{3}^{k_{3}},\;\mbox{ for }k_{1}+k_{2}+k_{3}=k$
as a linear combination of function of the form $(x_{1}+lx_{2}+jx_{3})^{k}$, with $j=0,1,\ldots,k$, $l=0,1,\ldots,k-k_{3}$.
For any $N>3$, we continue this process. Any monomial of the form
$x_{1}^{k_{1}}x_{2}^{k_{2}}\cdots x_{N}^{k_{N}},\;\;k_{1}+k_{2}+\ldots,k_{N}=k$
can be written as a linear combination of functions of the form
$(a_{1}x_{1}+a_{2}x_{2}+\ldots+a_{N}x_{N})^{k}$
with $a_{1}=1$; $a_{2}=0,\ldots,k_{1}$; $a_{3}=0,\ldots,k_{1}+k_{2}$; $\ldots$; and $a_{n}=0,1,\ldots,k$.
This proves the theorem. ∎