### v2

# Cauchy activation function and XNet

Abstract
We have developed a novel activation function, named the Cauchy Activation Function. This function is derived from the Cauchy Integral Theorem in complex analysis and is specifically tailored for problems requiring high precision. This innovation has led to the creation of a new class of neural networks, which we call (Comple)XNet, or simply XNet.
We will demonstrate that XNet is particularly effective for high-dimensional challenges such as image classification and solving Partial Differential Equations (PDEs). Our evaluations show that XNet significantly outperforms established benchmarks like MNIST and CIFAR-10 in computer vision, and offers substantial advantages over Physics-Informed Neural Networks (PINNs) in both low-dimensional and high-dimensional PDE scenarios.
1 Introduction
In today’s scientific exploration, the rise of computational technology has marked a significant turning point. Traditional methods of theory and experimentation are now complemented by advanced computational tools that tackle the complexity of real-world systems. Machine learning, particularly deep neural networks, has led to breakthroughs in fields like image processing and language understanding [3, 7], and its application to scientific problems–such as predicting protein structures [11, 12] or forecasting weather [15]–demonstrates its potential to revolutionize our approach.
One of the primary challenges in computational mathematics and artificial intelligence (AI) lies in determining the most appropriate function to accurately model a given dataset. In machine learning, the objective is to leverage such functions for predictive purposes. Traditional methods rely on predetermined classes of functions, such as polynomials or Fourier series, which, though simple and computationally manageable, may limit the flexibility and accuracy of the fit. In contrast, modern deep learning neural networks primarily employ locally linear functions with nonlinear activations.
While the trend in deep learning has been towards increasingly deeper architectures (from 8-layer AlexNet to 152/1001-layer ResNets), our work demonstrates an alternative approach. The Cauchy activation function’s superior approximation capabilities allow us to achieve comparable or better performance with significantly simplified architectures. For example, in our MNIST experiments, we reduced three fully connected layers to a single layer while maintaining high accuracy. This suggests that the power of neural networks may not solely lie in depth, but also in the choice of activation functions that can capture complex patterns more efficiently.
This finding has important implications for both theoretical understanding and practical applications:
1. Computational Efficiency: Fewer layers mean reduced computational costs and memory requirements
2. Training Stability: Shallower networks are typically easier to train and less prone to vanishing gradient problems
3. Interpretability: Simpler architectures may be more interpretable than very deep networks
1.1 Algorithm Development
In our previous work [1], we introduced the initial concept of extending real-valued functions into the complex domain, using the Cauchy integral formula to device a machine learning algorithm, and demonstrating its high accuracy and performance through examples in time series. In this work, we introduce a more general method stemming from the same mathematical principles. This approach is not confined to addressing mathematical physics problems such as low-dimensional PDEs; it also effectively tackles a broad spectrum of AI application issues, including image processing. This paper primarily showcases examples in image processing and both low-dimensional and high-dimensional (100-dimensional) PDEs. We are actively continuing our research to explore further applications, with very promising preliminary results.
Recent advancements in deep learning for solving PDEs and CV capabilities are well-documented, with significant contributions from transformative architectures that integrate neural networks with PDEs to enhance function approximation in high-dimensional spaces [23, 24, 25, 26, 29]. In CV, comprehensive surveys and innovative methods have significantly advanced visual processing techniques [30, 31, 32, 33].
Also, we have extensively reviewed literature on integrating complex domains into neural network architectures. Explorations of complex-valued neural networks and further developments highlight the potential of complex domain methodologies in modern neural network frameworks [2, 13, 14, 39, 46, 47].
Feedforward Neural Networks (FNNs), despite their capabilities, are often limited by the granularity of approximation they can achieve due to traditional activation functions such as ReLU or Sigmoid [40, 41, 4]. Recent works aim to optimize network performance through the discovery of more effective activation functions, leading to significant advancements in network functionality and computational efficiency [48, 16, 17, 18, 19, 20, 21, 34]. Adaptive activation functions have demonstrated significant improvements in convergence rates and accuracy, particularly in deep learning and physics-informed neural networks [35, 36, 37]. A recent survey [38] highlights their critical role in enhancing neural network performance across various tasks.
We propose a novel method that does not directly incorporate complex numbers. Drawing on insights from the complex Cauchy integral formula, we utilize Cauchy kernels as basis functions and introduce a new Cauchy activation function. This innovative approach significantly enhances the network’s precision and efficacy across various tasks, particularly in high-dimensional spaces.
Figure 1: ReLU
Figure 2: Sigmoid
The Cauchy activation function can be expressed as follows
$\phi_{\lambda_{1},\lambda_{2},d}(x)=\frac{\lambda_{1}*x}{x^{2}+d^{2}}+\frac{\lambda_{2}}{x^{2}+d^{2}},$
where $\lambda_{1},\lambda_{2},d$ are trainable parameters.
Theoretically, the Cauchy activation function can approximate any smooth function to its highest possible order. Moreover, from the figure below, we observe that our activation function is localized and decays at the both ends. This feature turned out to be very useful in approximating local data. This capability to finely tune to specific data segments sets it apart significantly from traditional activation functions such as ReLU.
Figure 3: 2 terms of Cauchy activation, with $\lambda_{1}=\lambda_{2}=d=1$
In Section 2, we will explain how the Cauchy activation function is derived and why it is mathematically efficient.
1.2 Enhanced Neural Network Efficiency with Cauchy Activation Function: High-Order Approximation and Beyond
Recent advancements in machine learning have been pivotal in tackling complex scientific challenges, such as high-dimensional Partial Differential Equations (PDEs) and nonlinear systems [23, 24, 26]. Traditional neural networks that employ activation functions like ReLU or Sigmoid often necessitate deep, multi-layered architectures to adequately approximate complex functions. This requirement typically results in increased computational complexity and extensive parameter demands.
In contrast, our innovative approach utilizes a single-layer network equipped with the Cauchy activation function, derived directly from our Cauchy Approximation Theorem (Theorem 1). This methodology introduces a fundamentally different and more efficient approximation mechanism:
$\left|f-\sum_{j=1}^{m}\frac{\lambda_{j}}{(\xi_{1}^{j}-z_{1})\cdots(\xi_{N}^{j}-z_{N})}\right|_{L^{\infty}(M)}\leq Cm^{-k},$ (1)
where $m$ represents the number of network parameters. This approximation exhibits an $O(m^{-k})$ convergence rate for any $k>0$, a substantial improvement over traditional ReLU networks, which typically achieve $O(m^{-r/d})$ convergence for $r$-smooth functions. Here, d represents the input dimensionality, and r denotes the smoothness degree of the function, as detailed in [27, 28].
Given a desired approximation error $\epsilon>0$, the comparative network sizes required to achieve this error are significantly reduced:
$m_{\text{ReLU}}=O\left((1/\epsilon)^{d/r}\right)\quad\text{versus}\quad m_{\text{Cauchy}}=O\left((1/\epsilon)^{1/k}\right),$ (2)
where $d$ and $r$ denote the input dimensionality and smoothness of the function, respectively, highlighting the drastic reduction in complexity that the Cauchy activation function enables.
This theoretical advantage has been translated directly into empirical performance improvements. Detailed examples and results of these experiments are discussed in Section 3.
In our MNIST experiments, even a simple two-layer network [128, 64] equipped with the Cauchy activation function outperformed a similar ReLU-based network, achieving a validation accuracy of 96.71% compared to 96.22%. Remarkably, a single-layer network [100] equipped with the Cauchy activation function achieved a competitive 96.30% accuracy while requiring significantly fewer epochs to converge compared to traditional activations.
The effectiveness of the Cauchy activation function extends across a range of tasks, from image classification, where it achieved 91.42% accuracy on CIFAR-10, outperforming ReLU’s 90.91%, to high-dimensional PDE solving, where it reduced error from 0.0349 to 0.00354.
2 Approximation Theorems
Consider, for example, a dataset comprising values of certain function $g(x_{i}),i=1,\ldots,n$ corresponding to points $x_{1},\ldots,x_{n}$ on the real line. These data may contain noises. We begin by assuming that the target function for fitting is a real-analytic function $f$. Although this assumption may appear stringent, it is important to note that non-analytic functions, due to their unpredictable nature and form, exhibit a weak dependence on specific data sets. Another perspective is that our aim is to identify the most suitable analytic function that best fits the provided dataset.
Real analytic functions can be extended to complex plane. The central idea of our algorithm is to place observers in the complex plane. Similar to activation for each node in artificial neural network, a weight is computed and assigned to each observer. The value of the predicted function $f$ at any point is then set to be certain weighted average of all observers. Our core mathematical theory is the Cauchy Approximation Theorem (Section 2, Theorem 1), derived in the next section. The Cauchy Approximation Theorem guarantees the efficacy and the accuracy of the predicted function. Comparing with the Universal Approximation Theorem for artificial neural network, whose proof takes considerable effort, our theorem comes directly from Cauchy Integration formula (eq 3).
In section 2, we start with Cauchy integral for complex analytical functions, leading us to the derivation the Cauchy Approximation Theorem. This theorem serves as the mathematical fundation of our algorithm. Theoretically, our algorithm can achieve a convergence rate of arbitrarily high order in any space dimensions.
In Sections 3, we evaluate our algorithm using a series of test cases. Observers are manually positioned within the space, and some known functions are employed to generate datasets in both one-dimensional and two-dimensional spaces. The predicted functions are then compared with the actual functions. As anticipated, the results are exceptional. We also conducted tests on datasets containing random noises, and the outcomes were quite satisfactory. The algorithm demonstrates impressive predictive capabilities when it processes half-sided data to generate a complete function.
We anticipate that this innovative algorithm will find extensive applications in fields such as computational mathematics, machine learning, and artificial intelligence. This paper provides a fundamental principle for the algorithm.
We will formally state the fundamental theory behind our algorithm.
Given a function $f$ holomorphic in a compact domain $U\subset\mathbb{C}^{N}$ in the complex $N$ dimensional space. For simplicity, we assume that $U$ has a product structure, i.e., $U=U_{1}\times U_{2}\times\ldots\times U_{n}$, where each $U_{k}$, $k=1,\ldots,N$ is a compact domain in the complex plane. Let $P$ be the surface defined by
$P=\partial U_{1}\times\partial U_{2}\times\ldots\times\partial U_{N}$
Then a multi-dimensional Cauchy integral formula for $U$ is given by:
[…]
for all $(z_{1},\ldots,z_{N})\in U$.
The magic of the Cauchy Integration Formula lies in its ability to determine the value of a function at any point using known values of the function. This concept is remarkably similar to the principles of machine learning!
The Cauchy integral can be approximated, to any prescribed precision, by a Riemann sum over a finite number of points on $P$. We can simplify the resulting Riemann sum in the following form:
$f(z_{1},z_{2},\ldots z_{N})\approx\sum_{k=1}^{m}\frac{\lambda_{k}}{(\xi_{1}^{k}-z_{1})(\xi_{2}^{k}-z_{2})\cdots(\xi_{N}^{k}-z_{N})},$ (4)
where $\lambda_{1},\lambda_{2},\ldots,\lambda_{m}$ are parameters depending on the functional values at the sample points on $P$.
The Cauchy integral Formula guarantees the accuracy of the above approximation if enough points on $P$ are taken. However, there is no reason at all that the sample points has to be on $P$ to achieve the best approximation. Indeed, the surface $P$ itself is quite arbitrary.
We can state our fundamental theorem.
Theorem 1 (Cauchy Approximation Theorem).
Let $f(z_{1},z_{2},\ldots z_{N})$ be an analytic function in an open domain $U\subset\mathbb{C}^{N}$ and let $M\subset U$ be a compact subset in $U$. Given any $\epsilon>0$, there is a list of points $(\xi_{1}^{k},\ldots,\xi_{N}^{k})$, for $k=1,2,\ldots,m$, in $U$ and corresponding parameters $\lambda_{1},\lambda_{2},\ldots,\lambda_{m}$, such that
$\vline\,f(z_{1},z_{2},\ldots z_{N})-\sum_{k=1}^{m}\frac{\lambda_{k}}{(\xi_{1}^{k}-z_{1})(\xi_{2}^{k}-z_{2})\cdots(\xi_{N}^{k}-z_{N})}\,\vline<\epsilon,$ (5)
for all points $(z_{1},z_{2},\ldots z_{N})\in M$.
As we have explained, the proof is a simple application of the Cauchy Theorem. We omit the details of the proof. We remark that, for any given $\epsilon>0$, the number of points $m$ needed is approximately at the level of $m\sim\epsilon^{-N}$, or equivalently, the error is approximately at $\epsilon\sim m^{-1/N}$. In fact, due to the nature of complex analytic functions, where the size of the function also bounds the derivative of the function, the error is much smaller. In fact, for any fixed integer $k>0$, one can show in theory that $\epsilon\sim o(m^{-k})$ for large $m$.
The Cauchy approximation is stated for complex analytical functions. If the original function is real, it can be simplified as follows.
$f(x_{1},x_{2},\ldots x_{N})\approx\mbox{Re}\left(\sum_{k=1}^{m}\frac{\lambda_{k}}{(\xi_{1}^{k}-x_{1})(\xi_{2}^{k}-x_{2})\cdots(\xi_{N}^{k}-x_{N})}\right),$ (6)
where $x_{1},x_{2},\ldots,x_{N}$ are real, $\lambda_{k},\;k=1,\ldots,m$ and $\xi_{i}^{j},\;i=1,\ldots N,\;j=1,\ldots,m$ are complex numbers with positive imaginary parts, Re($z$) is the real part of $z$.
A special simple case is when $N=1$, a function of one complex or real variable. In this case, the approximation is
$f(x)\approx\mbox{Re}\left(\sum_{k=1}^{m}\frac{\lambda_{k}}{\xi_{k}-x}\right),$ (7)
with $x$ real and $\xi_{k},\lambda_{k},\;k=1,\ldots m$ complex.
To derive the activation function, we express the complex parameters $\xi_{k}$ and weights $\lambda_{k}$ in terms of their real and imaginary parts:
$\xi_{k}=\xi_{k,\text{real}}+i\cdot\xi_{k,\text{imag}},\quad\lambda_{k}=\lambda_{k,\text{real}}+i\cdot\lambda_{k,\text{imag}}.$
The denominator $|\xi_{k}-x|^{2}$ can be written as:
$|\xi_{k}-x|^{2}=(\xi_{k,\text{real}}-x)^{2}+(\xi_{k,\text{imag}})^{2}.$
The fraction $\frac{\lambda_{k}}{\xi_{k}-x}$ is expanded as:
$\frac{\lambda_{k}}{\xi_{k}-x}=\frac{\lambda_{k,\text{real}}+i\cdot\lambda_{k,\text{imag}}}{(\xi_{k,\text{real}}-x)+i\cdot\xi_{k,\text{imag}}}.$
Taking the real part, we obtain:
$\text{Re}\left(\frac{\lambda_{k}}{\xi_{k}-x}\right)=\frac{\lambda_{k,\text{real}}\cdot(\xi_{k,\text{real}}-x)+\lambda_{k,\text{imag}}\cdot\xi_{k,\text{imag}}}{(\xi_{k,\text{real}}-x)^{2}+(\xi_{k,\text{imag}})^{2}}.$
To simplify further, assume that the real and imaginary parts of $\lambda_{k}$ are trainable parameters, denoted as $\lambda_{1}$ and $\lambda_{2}$, and let $\xi_{k,\text{imag}}=d$, where $d$ is another trainable parameter. Substituting these into the expression, we obtain the activation function:
$\phi_{\lambda_{1},\lambda_{2},d}(x)=\frac{\lambda_{1}\cdot x}{x^{2}+d^{2}}+\frac{\lambda_{2}}{x^{2}+d^{2}}.$
Explanation of Trainable Parameters:
  1. 1.
    $\lambda_{1}$: This controls the linear contribution of the input $x$. Every function can be decomposed into odd and even parts, this term captures the odd part.
  2. 2.
    $\lambda_{2}$: This adjusts the constant term, providing additional flexibility in shaping the activation. This term captures the even part of the function.
  3. 3.
    $d$: This parameter defines the scale of the denominator, controlling the range and smoothness of the activation function.
Figure 4: Visualization of the two terms of the Cauchy activation function under different parameter settings.
The Cauchy activation function provides unique advantages due to its flexibility and localization properties, making it particularly suitable for tasks involving smooth or highly localized data patterns. For effective initialization, we typically set $\lambda_{1}$ and $\lambda_{2}$ to small positive values (e.g., around 0.01), which ensures minimal bias and avoids large gradients during early training. The parameter $d$ is initialized to 1, providing a balance between smoothness and localization. These initial settings allow the activation function to stabilize training dynamics and adaptively learn optimized parameters to enhance model performance.
Cauchy approximation theorem can be easily implemented into an artificial neural network ([1]). The resulting network, termed CauchyNet, is very efficient for lower dimensional problems, say for $N\leq 10$. However, for large $m$, the multiplicative terms in the denominator pose serious computational difficulties. As typical computer vision problems are high dimensional, with the input data for a $30\times 30$ pixel image being at least 900 dimensional, we need a different algorithm to handle high dimensional problems.
Towards this end, we prove a general approximation theorem. The central idea is that, if one can approximate one-dimensional function with certain functional class, then one can extend the functional class to approximate functions in any dimension. Mathematically, this corresponds to approximation in dual spaces. This method is particularly effective for feature-capturing.
As an example, Cauchy approximation can approximate any one dimensional function, as we see in the above theorem, therefore it can approximate any dimensioanl function through linear combinations.
More precisely, we consider $C(\mathbb{R})=C(\mathbb{R},\mathbb{R})$, the set of continuous real-valued functions in $\mathbb{R}$. A family of real valued functions on $\mathbb{R}$, $\Phi=\{\phi_{a}\;|\;a\in A\}$, where $A$ is some index set, is said to have universal approximation property or simply approximation property, if for any closed bounded interval $I$ in $\mathbb{R}$ and any continuous function $g\in C(I,\mathbb{R})$, there is a sequence of functions $g^{j}=a_{1}^{j}g_{a_{1}}+a_{2}^{j}g_{a_{2}}+\dots+a_{k_{j}}^{j}g_{a_{k_{j}}}$ such that $g^{j}\rightarrow g$ uniformly over $I$. In other words, $\Phi$ has approximation property if every continuous function can be approximated by a linear combination of functions in $\Phi$, uniformly over a bounded interval.
The main result is the following general approximation theorem.
Theorem 2 (General Approximation Theorem).
Let $\Phi$ be a family of functions in $C(\mathbb{R},\mathbb{R})$ with the universal approximation property. Let
$\Phi^{N}=\{\phi_{a}(a_{1}x_{1}+a_{2}x_{2}+\ldots+a_{N}x_{N})\;|\;(a_{1},\ldots,a_{N})\in\mathbb{R}^{N},\phi_{a}\in\Phi\}$
Then, the family of functions $\Phi^{N}$ has the universal approximation property in $C(\mathbb{R}^{N},\mathbb{R})$, i.e, every continuous function in $\mathbb{R}^{N}$ can be approximated by a linear combination of functions in $\Phi^{N}$, uniformly over a compact subset in $\mathbb{R}^{N}$.
The proof of this theorem can be found in Appendix A.
The Cauchy Approximation Theorem and the General Approximation Theorem demonstrate that the Cauchy activation function can be used effectively, at least mathematically, to approximate functions of any dimension. We will test its practical applications in the following sections.
Compared to popular neural networks, it appears that only the activation function needs to be changed. However, due to the novelty of the Cauchy activation function, structural changes to the neural networks are necessary to leverage its intrinsic efficiencies. Generally, we can significantly simplify complex networks to achieve comparable or better results. This theoretical and operational framework prompts us to rename the network to ’XNet’ to reflect the fundamental improvements in our networks capabilities.
3 Examples
In this section, we focus on image classification and PDEs. Regression tasks, especially function fitting and computations for other PDEs, are also highly relevant and can be explored further, as detailed in [8]. These tasks serve as important benchmarks for evaluating the approximation capabilities of neural networks and their ability to generalize under different problem settings.
Our approach was primarily compared with MLP, where it demonstrated significant advantages in performance. Specifically, the Cauchy activation function showcased superior approximation power in regression tasks by achieving high-order accuracy in function fitting. This is particularly evident in its ability to handle complex target functions with minimal error, as well as its robustness in noisy data scenarios.
Moreover, when applied to PDEs, the method consistently achieved higher accuracy, usually much more than 10 times while achieved in much less computing time, in solving both linear and nonlinear equations, underscoring its potential in scientific computing applications. These results demonstrate the versatility of our approach across diverse tasks, and future work will further investigate its applicability in more challenging domains.
It is important to note that Theorem 2 applies to any higher orders. When solving PDEs, the higher-order accuracy of the solution is a key criterion for evaluating the method. As we optimize through parameter tuning, we will observe increasingly higher-order approximation effects. With the amount of data we have, we have already seen high-order accuracy. Since our theorem applies to arbitrary higher orders, we can observe even higher-order accuracy through parameter tuning and other techniques. This will be a focus of our future work.
3.1 Regression Task
Our approach utilizes the Cauchy activation function known for its ability to achieve high-order accuracy in function fitting.
3.1.1 High-Order Approximation Analysis
In a noise-free environment, we define the target function for the regression task as:
$y_{\text{train}}=x_{1}^{2}-x_{1}\cdot x_{2}+3\cdot x_{2}+x_{2}^{2}+\frac{1}{5+x_{1}^{2}}.$
We conducted training with a dataset size of $N=2500$, using a single hidden layer with 400 neurons, over 12000 epochs, and a learning rate of 0.001. The observed MSE was remarkably low at $1.7\times 10^{-6}$, indicating a high level of model accuracy.
The error dynamics for the Cauchy model are given by:
$\text{MSE}=O\left(\frac{1}{N}+\frac{1}{h^{p}}\right),$
where $h=400$ is the number of neurons, and $p$ represents the order of approximation. Solving for $p$, we find:
$1.7\times 10^{-6}=\frac{1}{2500}+\frac{1}{400^{p}},$
resulting in $p\approx 12.8$, a clear demonstration of high-order approximation capability.
3.1.2 Comparison with ReLU under Noisy Conditions
To simulate real-world scenarios in our regression tasks, we introduced Gaussian noise to the target function:
$y_{\text{train}}=x_{1}^{2}-x_{1}\cdot x_{2}+3\cdot x_{2}+x_{2}^{2}+\frac{1}{5+x_{1}^{2}}+\mathcal{N}(0,\sigma^{2}),$
where $\mathcal{N}(0,\sigma^{2})$ denotes Gaussian noise with a standard deviation of $\sigma=0.1$, representing a low noise scenario.
In a comparative study over 1000 epochs, we assessed the performance of neural networks equipped with Cauchy and ReLU activation functions. We used a fixed learning rate of $lr=0.001$ across all experiments. Although we tested various learning rates, $lr=0.001$ was chosen as it allowed us to clearly discern performance differences between the activation functions within 1000 epochs.
Table 1: Loss Comparison: Cauchy vs. ReLU
Activation Hidden Dim Noise Level Loss
Cauchy 400 Clean 0.000538
Cauchy 400 10% Noise 0.010413
ReLU 400 Clean 0.010158
ReLU 400 10% Noise 0.021373
Cauchy 800 Clean 0.000258
Cauchy 800 10% Noise 0.010155
ReLU 800 Clean 0.001646
ReLU 800 10% Noise 0.012083
As we noted eralier, if we run more epoches with smaller learning rate, MSE with Cauchy activation can be further reduced to the order of $10^{-6}$, clearly showing the high order effect of Cauchy approximation. The data clearly demonstrate the superior performance of the Cauchy activation function, which maintained lower loss values across both noise levels and hidden dimensions, confirming its effectiveness and robustness in noisy conditions.
3.2 Handwriting Recognition: MNIST with XNet
The XNet architecture used for these experiments consists of an input layer $W_{in}$, a hidden layer $W$, and an output layer $W_{out}$. The input $28\times 28$ grayscale image is reshaped into a $784\times 1$ vector, which is compressed by the input layer into $N$ features and processed by the $N\times N$ hidden layer. The output layer maps these $N$ features to 10 classes representing digits 0–9. The model was implemented in PyTorch and trained using the ADAM optimizer with a learning rate of 0.0001.
We evaluated XNet on the MNIST dataset using two fully connected architectures–a single-layer network ([100]) and a two-layer network ([128, 64])–as well as a CNN with a simple multi-scale kernel size convolution layer. Across all architectures, the Cauchy activation function demonstrated superior performance in accuracy, F1 score, AUC, and other metrics on the testing data. These results underscore that the Cauchy activation outperforms other activation functions across multiple benchmarks, including accuracy, loss minimization, convergence speed, generalization error, F1 score, and AUC, albeit with a slightly longer runtime. The slightly longer runtime is due to the custom implementation of the Cauchy activation function compared to the optimized PyTorch-packaged activations.
In the [100] network, the Cauchy activation achieved a validation accuracy of $96.33\%$, a validation loss of $0.1379$, an F1 score of $0.964$, and an AUC of $0.987$, significantly outperforming ReLU ($95.27\%$, $0.1693$, $0.952$, $0.978$) and Sigmoid ($90.73\%$, $0.3289$, $0.907$, $0.940$). Similarly, in the [128, 64] network, Cauchy led with a validation accuracy of $96.87\%$, a validation loss of $0.1434$, an F1 score of $0.969$, and an AUC of $0.990$, surpassing ReLU ($96.22\%$, $0.1628$, $0.962$, $0.985$) and Leaky ReLU ($96.46\%$, $0.1547$, $0.965$, $0.987$). While the Cauchy activation incurred a slightly longer runtime due to our custom implementation compared to PyTorch’s built-in functions, its substantial improvements in accuracy, generalization, and other metrics validate its efficiency and robustness for diverse applications.
Figures 9, 14, 9, and 14 depict the training and validation loss curves with learning rate 0.001 (results for learning rate 0.01 show similar patterns and are omitted for brevity). The Cauchy activation (pink curve) demonstrates notably faster convergence during training, not only achieving near-zero training loss more quickly but also reaching optimal validation metrics in earlier epochs. This rapid convergence to optimal performance is particularly valuable in practical applications. Sigmoid (orange curve), on the other hand, converges slowly and suffers from high training and validation losses due to vanishing gradients. While ReLU and Leaky ReLU show competitive final performance, Cauchy outperforms them in terms of convergence speed and the ability to achieve better metrics earlier in the training process.
Cauchy’s efficiency is further highlighted by its ability to achieve high validation accuracy, low validation losses, and superior F1 and AUC scores in earlier epochs compared to other activation functions. This characteristic makes it particularly well-suited for tasks requiring both high precision and efficient training, such as medical diagnosis or fraud detection. While the final epoch results show some signs of overfitting, the superior performance achieved in earlier epochs demonstrates Cauchy’s potential to reach optimal solutions more efficiently. Although Leaky ReLU achieves competitive validation loss in later epochs, it requires more training time to reach comparable performance levels. Sigmoid consistently exhibits the weakest generalization throughout the training process.
These results underscore the potential of the Cauchy activation function not only in achieving superior metrics but also in reaching optimal performance more efficiently, making it a promising choice for neural network architectures, particularly in scenarios requiring both high precision and training efficiency. The rapid convergence to optimal performance suggests that with proper regularization and early stopping strategies, Cauchy activation could provide both better results and more efficient training processes.
Figure 5: Training Loss for [100].
Figure 6: Validation Loss for [100].
Figure 7: F1 Score for [100].
Figure 8: AUC for [100].
Figure 9: Performance Metrics for the [100] Model.
Figure 10: Training Loss for [100].
Figure 11: Validation Loss for [128,64].
Figure 12: F1 Score for [128,64].
Figure 13: AUC for [128,64].
Figure 14: Performance Metrics for the [128,64] Model.
Params Epoch Train Loss Train Acc Val Loss Val Acc F1 Score AUC Gen Error Time (s)
$[100],0.001,\text{relu}$ 20 0.1337 0.9643 0.1651 0.9524 0.9522 0.9979 0.0120 1.08
$[100],0.001,\text{sigmoid}$ 20 0.3088 0.9165 0.3259 0.9079 0.9070 0.9923 0.0086 1.21
$[100],0.001,\text{tanh}$ 20 0.1938 0.9453 0.2224 0.9363 0.9361 0.9960 0.0090 1.09
$[100],0.001,\text{swish}$ 20 0.1984 0.9444 0.2279 0.9351 0.9348 0.9959 0.0093 1.23
$[100],0.001,\text{gelu}$ 20 0.1770 0.9513 0.2060 0.9397 0.9393 0.9966 0.0116 1.10
[…]
$[100],0.001,\text{cauchy}$ 20 0.0079 0.9982 0.1929 0.9630 0.9627 0.9986 0.0352 1.20
$[100],0.01,\text{relu}$ 20 0.0755 0.9785 0.1355 0.9594 0.9592 0.9989 0.0192 1.09
$[100],0.01,\text{sigmoid}$ 20 0.2137 0.9394 0.2382 0.9313 0.9308 0.9962 0.0082 1.15
$[100],0.01,\text{tanh}$ 20 0.0989 0.9705 0.1296 0.9606 0.9606 0.9988 0.0098 1.08
Table 2: Performance Comparison of Activation Functions (Part 1: $[100],0.001$ and $[100],0.01$)
Params Epoch Train Loss Train Acc Val Loss Val Acc F1 Score AUC Gen Error Time (s)
$[128,64],0.001,\text{relu}$ 20 0.0687 0.9811 0.1165 0.9663 0.9663 0.9989 0.0147 1.15
$[128,64],0.001,\text{sigmoid}$ 20 0.2849 0.9183 0.3011 0.9117 0.9106 0.9929 0.0066 1.15
$[128,64],0.001,\text{tanh}$ 20 0.1169 0.9666 0.1603 0.9532 0.9531 0.9978 0.0135 1.14
$[128,64],0.001,\text{swish}$ 20 0.1217 0.9652 0.1663 0.9513 0.9512 0.9978 0.0140 1.11
$[128,64],0.001,\text{gelu}$ 20 0.0990 0.9713 0.1454 0.9548 0.9546 0.9983 0.0166 1.11
[…]
$[128,64],0.001,\text{cauchy}$ 20 0.0632 0.9796 0.1141 0.9671 0.9671 0.9992 0.0125 1.35
$[128,64],0.01,\text{relu}$ 20 0.0503 0.9837 0.1066 0.9660 0.9656 0.9993 0.0176 1.12
$[128,64],0.01,\text{sigmoid}$ 20 0.1667 0.9499 0.1826 0.9460 0.9458 0.9978 0.0038 1.15
$[128,64],0.01,\text{tanh}$ 20 0.0682 0.9785 0.1207 0.9627 0.9622 0.9989 0.0158 1.14
[…]
$[128,64],0.01,\text{swish}$ 20 0.0650 0.9785 0.1157 0.9654 0.9653 0.9991 0.0131 1.25
$[128,64],0.01,\text{gelu}$ 20 0.0521 0.9824 0.0993 0.9678 0.9676 0.9993 0.0147 1.13
$[128,64],0.01,\text{cauchy}$ 20 0.0632 0.9796 0.1141 0.9671 0.9671 0.9992 0.0125 1.35
Table 3: Performance Comparison of Activation Functions (Part 2: $[128,64],0.001$ and $[128,64],0.01$)
As shown in Tables 1 and 2, the Cauchy activation function demonstrates superior performance across different network architectures and learning rates. With the [100] architecture and learning rate of 0.001, it achieves remarkable training accuracy of $99.82\%$ and validation accuracy of $96.30\%$, significantly outperforming other activation functions. Similarly impressive results are observed with the [128, 64] architecture, where Cauchy maintains consistently high performance with $97.96\%$ training accuracy and $96.71\%$ validation accuracy at learning rate 0.01. The higher generalization error observed (e.g., 0.0352 for [100] architecture) can be attributed to the rapid convergence and the absence of regularization techniques in our comparative setup, rather than an inherent limitation of the activation function. Notably, Cauchy activation achieves superior AUC scores (0.9986-0.9992) across all configurations, indicating excellent classification reliability. While the computational time appears longer in our experiments, this is primarily due to our custom implementation of the Cauchy activation function compared to PyTorch’s highly optimized built-in functions for other activations, rather than an inherent computational complexity of the function itself. A native implementation would likely achieve comparable computational efficiency. These results suggest that the Cauchy activation function, when properly regularized and efficiently implemented, could be the optimal choice among common activation functions for similar classification tasks.
In the 2nd experiment, we used a simple convolutional neural network (CNN) with a single convolutional layer, followed by a max-pooling layer and a fully connected layer.
Figure 15: Training and validation performance of CNN with ReLU and Cauchy activation functions.
We experimented with two different activation functions: ReLU and Cauchy. In our model configuration, we replaced the ReLU activation function with the Cauchy activation function after the third convolutional layer. The learning rate was set to 0.001 for both models. The training and validation losses, as well as accuracies after 20 epochs, are presented in Table 4.
Table 4: Performance comparison of CNN with ReLU and Cauchy activation functions after 20 epochs, using a 3-layer convolutional architecture.
Activation LR Training Loss Training Acc Valid Loss Valid Acc
ReLU 0.001 0.0190 0.9935 0.0312 0.9905
Cauchy 0.001 0.0039 0.9988 0.0254 0.9924
As shown in Table 4, the model using the Cauchy activation function achieved a training loss of 0.0039 and a training accuracy of $99.88\%$, while the validation loss was 0.0254 and the validation accuracy was $99.24\%$. On the other hand, the model using the ReLU activation function resulted in a training loss of 0.0190 and a training accuracy of $99.35\%$, with a validation loss of 0.0312 and a validation accuracy of $99.05\%$.
These results suggest that the Cauchy activation model outperformed the ReLU model in both training and validation, showing lower losses and higher accuracies, indicating better generalization capabilities.
Remark
The fluctuations observed in the validation loss curve for the Cauchy activation function are a result of its unique derivative properties. Unlike ReLU, whose derivative is piecewise constant, the derivative of the Cauchy activation function exhibits significant variations across the input space. This characteristic allows the Cauchy activation function to capture complex, nonlinear relationships more effectively, particularly in challenging regions of the loss landscape.
These fluctuations can be more pronounced when using a relatively larger learning rate, as the optimizer takes larger steps during training, amplifying the sensitivity to the Cauchy activation’s gradient dynamics. Importantly, reducing the learning rate mitigates this behavior, leading to smoother convergence. The trade-off lies in the balance between learning speed and stability: smaller learning rates improve stability, while larger rates accelerate convergence but may introduce short-term oscillations.
Despite these initial fluctuations, the validation accuracy curve shows stable improvement, indicating that the Cauchy activation function’s high-order approximation capability ultimately enables superior performance.
3.3 CIFAR-10
We conducted a comprehensive study to evaluate the performance of the proposed Cauchy activation function compared to six widely used activation functions: ReLU, Sigmoid, Tanh, Swish, GeLU, and Leaky ReLU. The experiments were performed on the CIFAR-10 dataset, which contains 60,000 32x32 color images across 10 classes, making it a standard benchmark for image classification tasks.
Experimental Setup:
  • CNN Architecture: A custom Convolutional Neural Network (CNN) with 6 convolutional layers followed by fully connected layers. For experiments with the proposed Cauchy activation function, the original three fully connected layers were replaced with a single fully connected layer. Additionally, a normalization block (NB) was introduced to enhance stability. The Cauchy activation was applied exclusively in this modified fully connected layer, leveraging its high-order approximation capability. For all other activation functions (ReLU, Sigmoid, Tanh, Swish, GeLU, Leaky ReLU), the architecture remained in its original form, with the activation functions directly replacing ReLU without altering the structure.
  • ResNet9 Architecture: A compact version of the ResNet architecture, designed for efficiency on smaller datasets. For experiments with the Cauchy activation function, the latter half of the convolutional layers and the residual connections were modified to use Cauchy activation. This adjustment aimed to explore the activation’s impact on deeper network components. For all other activation functions, ReLU was directly replaced with the corresponding activation function throughout the network, maintaining the original structure and residual connections.
Training Procedure: For both architectures, training was conducted in two phases:
  1. 1.
    An initial training phase with $20$ epochs using a learning rate of $0.001$.
  2. 2.
    A fine-tuning phase with $10$ epochs using a reduced learning rate of $0.0001$.
Results and Analysis:
Table 5: Comparison of Activation Functions on CIFAR-10 (6 Convolutional Layers, 30 Epochs)
Activation Function Final Validation Accuracy (%)
ReLU 78.60
Sigmoid 76.71
Tanh 76.74
Swish 78.54
GeLU 78.81
Leaky ReLU 79.29
Cauchy 81.90
As shown in Table 5, the Cauchy activation function achieves the highest accuracy of 81.90%, outperforming other activation functions such as ReLU (78.60%) and GeLU (78.81%).
Table 6: Performance Comparison of Activation Functions on CIFAR-10 with ResNet9
Activation Function Learning Rate Validation Accuracy (%)
ReLU 0.01 90.91
ReLU 0.005 90.72
Sigmoid 0.01 80.30
Sigmoid 0.005 80.86
Tanh 0.01 85.99
Tanh 0.005 84.84
Swish 0.01 90.34
Swish 0.005 90.62
GeLU 0.01 90.57
GeLU 0.005 91.09
Leaky ReLU 0.01 90.56
Leaky ReLU 0.005 90.65
Cauchy 0.01 90.28
Cauchy 0.005 91.42
In the ResNet9 experiments, summarized in Table 6, the Cauchy activation function achieves the highest accuracy (91.42%) at $\text{lr}=0.005$, outperforming GeLU (91.09%). Lower learning rates consistently improve performance across all activation functions, highlighting the benefits of stable convergence.
Conclusion: The results demonstrate the superiority of the Cauchy activation function in terms of accuracy and generalization. Its ability to achieve higher accuracy with fewer fully connected layers in CNN experiments further validates its theoretical high-order approximation capabilities. Moreover, its competitive performance in ResNet9 experiments suggests that Cauchy activation can be a promising alternative to standard activation functions in deep learning.
3.4 PDE: Heat Function
The 1-dimensional heat equation in this example is given by:
$\frac{\partial u}{\partial x}-2\frac{\partial u}{\partial t}-u=0,$ (8)
where $u(x,t)$ is the temperature distribution function, $x$ is the spatial coordinate, and $t$ is the time.
The boundary and initial conditions are specified as follows:
  • Initial condition:
    $u(x,0)=6e^{-3x},\quad\text{for}\;0\leq x\leq 2.$ (9)
  • Boundary conditions:
    $u(0,t)=0,\quad\text{for}\;0\leq t\leq 1,$ (10)
    $u(2,t)=0,\quad\text{for}\;0\leq t\leq 1.$ (11)
We trained the original PINN with the sigmoid activation function and then modified the network to use a Cauchy activation function instead of the sigmoid function. The training process was repeated with the new activation function to compare the performance.
Figure 16: loss curve by XNet and PINN
Table 7: Comparison with PINN
Activation Network Layers Training Loss Mean Error
Sigmoid 5 Layers FNN 0.0064 2e-3
Cauchy 5 Layers 0.0003 6e-5
3.5 Poisson Equation with Dirichelet Boundary Condition
We study the Poisson equation
$\nabla^{2}u(x,y)=f(x,y),\quad\text{for }(x,y)\in\Omega,$ (12)
with Dirichlet boundary conditions:
$u(x,y)=0,\quad\text{for }(x,y)\in\partial\Omega,$ (13)
where the source term $f(x,y)=-8\pi^{2}\sin(2\pi x)\sin(2\pi y)$. The ground truth solution is $u(x,y)={\rm sin}(2\pi x){\rm sin}(2\pi y)$.
The dataset size is 2000, 1000 interior points and 1000 boundary points.
In this simple low-dimensional equation, we firstly found that using the least-squares method in MATLAB effectively solves the problem. We worked with 1000 interior points and 1000 boundary points, placing 400 observation points, which correspond to the boundary points in the Cauchy integral formula in complex space. The method is equivalent with our CauchyNet in [1]. We defined the boundary as an ellipse. (If we optimize the boundary points, we can achieve even higher computational accuracy.) The results are shown below:
Figure 17: Cauchy activation
Figure 18: Analytic solution
Figure 19: Pointwise Difference between two solutions
The $L^{2}$ error is computed as follows: $\text{diff}=U_{\text{grid}}-F_{\text{grid}},$where $U_{\text{grid}}$ represents the predicted solution and $F_{\text{grid}}$ represents the analytical solution.
The $L^{2}$ error (mean squared error) is given by:
$L^{2}_{\text{error}}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(\text{diff}_{i})^{2}},$ (14)
where $n$ is the total number of grid points. The $L^{2}$ error is 0.0076886.
Next, we used a PINN model with a structure of size $[2,200,1]$, consisting of one hidden layer. The learning rate was set to 0.001 for the first 7000 epochs, and reduced to 0.0001 for the final 1000 epochs. The optimizer used was Adam. We compared the performance of two different activation functions: the tanh activation function and the Cauchy activation function over a total of 8000 training epochs.
Our loss function consists of two parts: the first part represents the Mean Squared Error (MSE) of the residuals from the equation, while the second part accounts for the MSE of the boundary conditions:
Loss $=\frac{1}{N_{\text{interior}}}\sum_{i=1}^{N_{\text{interior}}}\left(\frac{\partial^{2}u(x_{i})}{\partial x_{1}^{2}}+\frac{\partial^{2}u(x_{i})}{\partial x_{2}^{2}}-y_{\text{train},i}\right)^{2}+\frac{1}{N_{\text{boundary}}}\sum_{j=1}^{N_{\text{boundary}}}\left(u(x_{j})-u_{\text{boundary},j}\right)^{2}.$
Since the loss values were initially large and decreased significantly over time, the overall loss curve did not clearly highlight the differences between the methods. To address this, we plotted the loss curve for the final 1000 epochs separately, making it evident that our method (Cauchy activation) holds a clear advantage over the original PINN.
Figure 20: Overall loss curve for PINN using Tanh and Cauchy activations.
Figure 21: Zoomed-in loss curve for the final 1000 epochs.
Figure 22: Comparison of loss curves for PINN using Tanh and Cauchy activations. The Cauchy activation shows superior convergence speed and stability.
Activation Function Training loss
Tanh 0.0349
Cauchy 0.00354
Table 8: Comparison of activation functions and their respective training loss.
The results clearly demonstrate that our XNet model significantly outperforms the standard activation functions, as seen in the graph where the Cauchy activation function consistently achieves a lower loss, showing superior effectiveness.
In addition to the one-hidden-layer model, we also tested a PINN model with two hidden layers, with a structure of size $[2,20,20,1]$. The results for this two-hidden-layer model were similar to those of the one-hidden-layer model. Both models were trained using a learning rate of 0.001. After 8000 epochs, the tanh-based PINN achieved a training error of 0.119, while the XNet (Cauchy activation function) achieved a significantly lower training error of 0.0382.
3.5.1 Burger’s equation
To validate the performance of our proposed Cauchy activation function, we included the Burger equation in our experiments. The Burger equation, commonly used in fluid dynamics, is defined as:
$\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}=\nu\frac{\partial^{2}u}{\partial x^{2}},$
where $u(x,t)$ represents the velocity field, and $\nu$ is the viscosity coefficient.
In our implementation, we leveraged a Physics-Informed Neural Network (PINN) to solve the equation. The network approximates the solution $u(x,t)$ and enforces the equation’s physical constraints by minimizing the residual:
$f(x,t)=\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}-\nu\frac{\partial^{2}u}{\partial x^{2}}.$
Modifications for Experiments
We simplified the network architecture by reducing its depth from 10 layers to 5 layers compared to the original design. This adjustment improved computational efficiency while maintaining the capacity to accurately model the solution.
In addition to this, we replaced traditional activation functions, such as tanh, with our custom Cauchy activation function in all layers. This modification was intended to leverage the rapid convergence and robust learning characteristics of the Cauchy activation function.
Results
The results demonstrated a remarkable improvement in training efficiency:
  • With the tanh activation function, the network required 1000 epochs to reduce the training loss to zero.
  • With the Cauchy activation function, the training loss converged to zero within 20 epochs, showcasing the superior convergence speed of our approach.
Figure 23: Training loss comparison for Burger’s equation using tanh and Cauchy activation functions.
These findings validate that the Cauchy activation function is particularly well-suited for physics-informed neural networks, where rapid convergence and efficient training are critical.
3.5.2 High Dimensional PDE
In this section we test the XNet solver in the case of an 100-dimensional AllenCahn PDE with a cubic nonlinearity.
The Allen-Cahn equation is a reaction-diffusion equation that arises in physics, serving as a prototype for the modeling of phase separation and order-disorder transitions (see, e.g., [10]). This equation is defined as:
$\frac{\partial u}{\partial t}(t,x)+u(t,x)-[u(t,x)]^{3}+\left(\Delta_{x}u\right)(t,x)=0,$
with the solution $u$ satisfying for all $t\in[0,T)$, $x\in\mathbb{R}^{d}$:
$u(T,x)=g(x).$
Assume for all $s,t\in[0,T]$, $x,w\in\mathbb{R}^{d}$, $y\in\mathbb{R}$, $z\in\mathbb{R}^{1\times d}$, $m\in\mathbb{N}$ that $d=100$, $T=\frac{3}{10}$, $\mu(t,x)=0$, $\sigma(t,x)w=\sqrt{2}w$, and $\Upsilon(s,t,x,w)=x+\sqrt{2}w$. The reaction term is defined as
$f(t,x,y,z)=y-y^{3},$
capturing the double-well potential of the Allen-Cahn equation, where the two minima at $y=-1$ and $y=1$ represent stable equilibrium states.
The terminal condition
$g(x)=\left[2+\frac{2}{5}\|x\|_{\mathbb{R}^{d}}^{2}\right]^{-1}$
ensures smooth decay for large $\|x\|$, aligning with the expected physical behavior in bounded domains. These assumptions simplify the Allen-Cahn equation and provide a well-posed high-dimensional problem to evaluate the performance of the proposed method.
In our study, we employed a model identical to the one discussed in the paper [9]. We simplified the original model by reducing the multilayer perceptron (MLP) component to a single layer, effectively halving the parameter count. Then, we substituted the activation function and evaluated performance differences. We set up an Adam optimizer with a learning rate of 0.005. The comparison model configuration remained the same as described in the original paper [9] to ensure a fair evaluation.
Table 9: Comparison of Training Loss between Cauchy and ReLU
Step Cauchy Loss ReLU Loss Factor of Reduction
0 $1.5698\times 10^{-1}$ $1.5637\times 10^{-1}$ $\approx 1.00$
100 $2.9323\times 10^{-3}$ $9.8792\times 10^{-2}$ $\approx 33.70$
200 $3.3228\times 10^{-3}$ $7.8492\times 10^{-2}$ $\approx 23.63$
500 $2.9574\times 10^{-3}$ $3.7295\times 10^{-2}$ $\approx 12.61$
1000 $2.9667\times 10^{-3}$ $1.1308\times 10^{-2}$ $\approx 3.81$
2000 $2.3857\times 10^{-3}$ $4.9802\times 10^{-3}$ $\approx 2.09$
Figure 24: loss curve of Allen Cahn
Appendix A Proofs for Theorem 2
Proof: The proof is based on Stone-Weierstrauss Theorem: the set of all polynomials is dense in $C(B,\mathbb{R})$ for any compact subset $B\subset\mathbb{R}^{N}$. Since any polynomial is a linear combination of monomials, we just need to show that linear combinations of functions in $\Phi^{N}$ can approximate any monomial.
Since $\Phi$ has the approximation property, it therefore can approximate $x^{k}$ for any integer $k$, this implies that $\Phi^{N}$ can approximate any function of the form
$(a_{1}x_{1}+a_{2}x_{2}+\ldots+a_{N}x_{N})^{k},\;\mbox{ for }k\in\mathbb{N},\;(a_{1},\ldots,a_{N})\in\mathbb{R}^{N}.$
For any fixed integer $k$, the multinomial expansion of the above function is a linear combination of $C(k+N-1,N-1)$ monomial terms such as
$x_{1}^{k_{1}}x_{2}^{k_{2}}\cdots x_{N}^{k_{N}},\;\;k_{1}+k_{2}+\ldots,k_{N}=k.$
Here $C(k+N-1,N-1)$ is combinatorial number $k+N-1$ choosing $N-1$. We claim that the converse is also true, i.e., every such monomial can be written as a linear combination of functions of the form
$(a_{1}x_{1}+a_{2}x_{2}+\ldots+a_{N}x_{N})^{k}.$
Our theorem will directly follow from this claim.
We prove the claim by induction on $N$.
For $N=1$, this is obviously true.
For $N=2$,
$(x_{1}+0x_{2})^{k}$ $=$ $x_{1}^{k}$
$(x_{1}+1x_{2})^{k}$ $=$ $x_{1}^{k}+C(k,1)x_{1}^{k-1}x_{2}+\ldots+C(k,k-1)x_{1}x_{2}^{k-1}+x_{2}^{k}$
$(x_{1}+2x_{2})^{k}$ $=$ $x_{1}^{k}+C(k,1)2x_{1}^{k-1}x_{2}+\ldots+C(k,k-1)2^{k-1}x_{1}x_{2}^{k-1}+2^{k}x_{2}^{k}$
$(x_{1}+3x_{2})^{k}$ $=$ $x_{1}^{k}+C(k,1)3x_{1}^{k-1}x_{2}+\ldots+C(k,k-1)3^{k-1}x_{1}x_{2}^{k-1}+3^{k}x_{2}^{k}$
$(x_{1}+kx_{2})^{k}$ $=$ $x_{1}^{k}+C(k,1)kx_{1}^{k-1}x_{2}+\ldots+C(k,k-1)k^{k-1}x_{1}x_{2}^{k-1}+k^{k}x_{2}^{k}$
This is a system of equations for monomials of the form $x_{1}^{j}x_{2}^{k-j}$. The coefficient matrix is obviously nonsingular, therefore each $x_{1}^{j}x_{2}^{k-j}$ can be solved as a linear combination of $(x_{1}+jx_{2})^{k}$, with $j=0,1,\ldots,k$.
For $N=3$, fix any real number $l$, we can expand $((x_{1}+lx_{2})+jx_{3})^{k}$ for $j=0,1,\ldots,k$ into the product of the form
$(x_{1}+lx_{2})^{k-k_{3}}x_{3}^{k_{3}},\;\mbox{ for }k_{3}=1,...,k-1.$
Same as the previous case with $N=2$, each of the above terms can be expressed as a linear combination of
$((x_{1}+lx_{2})+jx_{3})^{k},\;\mbox{ for }\;j=0,1,\ldots,k.$
Now for any fixed $k_{3}$, we choose $l=0,1,\ldots,k-k_{3}$ and expand
$(x_{1}+lx_{2})^{k-k_{3}}x_{3}^{k_{3}}.$
Similar to the case with $N=2$, we now obtain $k-k_{3}$ independent equations. From these equations, we can solve each monomial
$x_{1}^{k_{1}}x_{2}^{k_{2}}x_{3}^{k_{3}},\;\mbox{ for }k_{1}+k_{2}+k_{3}=k$
as a linear combination of function of the form $(x_{1}+lx_{2}+jx_{3})^{k}$, with $j=0,1,\ldots,k$, $l=0,1,\ldots,k-k_{3}$.
For any $N>3$, we continue this process. Any monomial of the form
$x_{1}^{k_{1}}x_{2}^{k_{2}}\cdots x_{N}^{k_{N}},\;\;k_{1}+k_{2}+\ldots,k_{N}=k$
can be written as a linear combination of functions of the form
$(a_{1}x_{1}+a_{2}x_{2}+\ldots+a_{N}x_{N})^{k}$
with $a_{1}=1$; $a_{2}=0,\ldots,k_{1}$; $a_{3}=0,\ldots,k_{1}+k_{2}$; $\ldots$; and $a_{n}=0,1,\ldots,k$.
This proves the theorem. ∎